Potassium is mildly radioactive because one of its natural isotopes is radioactive: 40K undergoes β- decay with a half-life of t1/2 = 1.248 × 10 9 years. The isotopic abundances for potassium are:    39K 93.258 % (atom percent) stable   

   40K 0.0117 at % β- decay

41K 6.730 at % stable

Calculate the activity (measured in curies or becquerel) of 1 cubic centimeter of KCl.
Answer #1

For a given mass m (in grams) of an isotope with atomic mass {\displaystyle m_{\text{a}}} (in g/mol) and a half-life of t_{1/2} (in s), the radioactivity can be calculated using:

{\displaystyle A_{\text{Bq}}={\frac {m}{m_{\text{a}}}}N_{\text{A}}{\frac {\ln 2}{t_{1/2}}}}------Eq(1)

With N_{{\text{A}}} = 6.02214076×1023 mol−1, the Avogadro constant.

Since {\displaystyle m/m_{\text{a}}} is the number of moles n , the amount of radioactivity A can be calculated by:

{\displaystyle A_{\text{Bq}}=nN_{\text{A}}{\frac {\ln 2}{t_{1/2}}}}------Eq(2)

Here given that, only 40K  undergoes Beta decay and it has 0.000117 atomic percent and the atomic mass of the potassium is 40 g/mol and the t1/2 given as 1.248 *109 years,

substuting all in the above eqaution 2,

ABq = 29.32 Bq.

Thank You